OpenStudy (anonymous):

Which of the points is a solution to -5x – 3y = 15?A. (-5, 0) B. (-3, 0) C. (1, 3) D. (3, 0)

OpenStudy (amistre64):

which one fits?

OpenStudy (angela210793):

subtitude every point in the equation and see which one fits

OpenStudy (amistre64):

this is to test your ability to understand what the notations represent

OpenStudy (amistre64):

wihtout knowing the notation, your pretty much not going to be able to handle anything based on the notation

OpenStudy (anonymous):

@amistre64: It's actually to "test 'our' ability to understand what the notations represent" :P

OpenStudy (amistre64):

...semantics lol

OpenStudy (anonymous):

I meant in the sense that OP is testing us lol

OpenStudy (amistre64):

this is the part of studying when you get to interact and be a part of the answer ... yay!!

OpenStudy (anonymous):

would it be B

OpenStudy (amistre64):

lets see :) -5x – 3y = 15 -3 0 ------------- 15 + 0 = 15 yes :)

OpenStudy (anonymous):

Okay Amistre, now we know that you can substitute, yay!

OpenStudy (amistre64):

its the acme of my mathing abilities ;)

OpenStudy (angela210793):

lolz xD

OpenStudy (anonymous):

lol :D